OpenStudy (x__hazel__x):

PLEASE HELP! 8. There are 10 marbles in a bag. Tyrell draws one marble out at a time without looking and replaces it in the bag. Tyrell did this 100 times. Tyrell drew red marbles 9 times, green marbles 53 times, and blue marbles 38 times. Based on this Tyrell's experiment, predict how many of each color marble is in the bag. A. 3 red, 3 green, and 4 blue B. 1 red, 5 green, and 4 blue C. 1 red,1 green, and 4 blue D. 3 red, 4 green, and 3 blue

OpenStudy (x__hazel__x):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

`Tyrell drew red marbles 9 times...Tyrell did this 100 times` there are 100 total trials. Nine of them result in red so the empirical probability of getting red is 9/100 = 0.09 multiply this with the number of marbles in the bag to get 0.09*10 = 0.9 which will round to 1 red marble in the bag

jimthompson5910 (jim_thompson5910):

`Tyrell drew...green marbles 53 times` we have 53 green out of 100 total what is the empirical probability of getting green?

OpenStudy (x__hazel__x):

I think 53/100 = 0.53 0.53 * 10 = 5.3

jimthompson5910 (jim_thompson5910):

good, that will round to 5 green marbles

OpenStudy (x__hazel__x):

Thank you!

jimthompson5910 (jim_thompson5910):

no problem